When a polynomial is divided by x + 1, the remainder is 5. When the same polynomial is divided by x – 1, the remainder is 1. When the same polynomial is divided by x – 2, the remainder is 20. Determine the remainder when the polynomial is divided (x + 1)(x – 1)(x – 2).


1
Expert's answer
2021-11-29T02:29:23-0500

Let "q_1(x), q_2(x), q_3(x)" be the quotient when polynomial f(x) is divided by (x+1), (x-1) and (x-2) respectively.

By Division Algorithm

Dividend = Divisor "\times" Quotient + Remainder

"f(x) = (x+1)q_1(x) + 5 \;\;\;(1) \\

f(x) = (x-1)q_2(x) +1 \;\;\;(2) \\

f(x) = (x-2)q_3(x) +20 \;\;\;(3)"

Put x = -1 in equation (1), x=1 in equation (2) and x = 2 in equation (3), we get

"f(-1) = 5 \\

f(1) = 1 \\

f(2) = 20"

Let "ax^2 +bx +c" be the remainder when polynomial f(x) is divided by (x+1)(x-1)(x-2) and quotient = q_n(x)

"f(x) = (x+1)(x-1)(x-2)q_n(x) + (ax^2 +bx+c) \\

f(-1) = a-b+c \\

a-b+c = 5 \;\;\;(4) \\

f(1) = a+b+c \\

a+b+c=1 \;\;\;(5) \\

f(2) = 4a+2b+c \\

4a+2b+c = 20 \;\;\;(6)"

Adding equation (4) and (5), we get

"2a+2c= 6 \\

a+c = 3 \;\;\;(7)"

Multiplying equation (5) by 2 and subtracting it from equation (6)

"4a+2b+c = 20 \\

2a+2b+2c = 2 \\

2a-c = 18"

Adding (7) and (8)

"a+c = 3 \\

2a-c = 18 \\

3a=21 \\

a = 7"

Using a in equation (7), c = -4

Using a and c in equation (5), b = -2

Remainder is "7x^2 -2x -4"